For the formation of hydrazine (ΔΗ n n -187.78 kJ and ASon -226.3 J/K), what is ΔG at l 000 K? Report your answer to 4 significant figures.) G 38.52 kJ At what temperature (in Kelvin) do the forward and reverse reactions occur in equilibrium? (Report your answer to 4 significant figures.) #K 丼Please show all work and formulas. Thanks
Answer #1

1)

ΔHo = -187.78 KJ

ΔSo = -226.3 J/K

= -0.2263 KJ/K

T = 1000 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -187.78 - 1000.0 * -0.2263

ΔGo = 38.52 KJ

Answer: 38.52 KJ

2)

For equilibrium, ΔG = 0

ΔG = 0.0 KJ

ΔHo = -187.78 KJ

ΔSo = -226.3 J/K

= -0.2263 KJ/K

use:

ΔG = ΔH - T*ΔS

0.0 = -187.78 - T *-0.2263

T = 830. K

Answer: 830. K